OpenStudy (sleepyjess):

How many moles of calcium hydroxide are needed to produce 73.7 milliliters of water, if the density of water is 0.987 g/mL? HCl + Ca(OH) 2 ---> H2O + CaCl2 @Abhisar

OpenStudy (sleepyjess):

2 HCl + Ca(OH)2 → 2 H2O + CaCl2 Mass of water: m=d*v =.987 * 73.7 = 72.74 Molar mass of water is Hydrogen ((1.008) * 2) + Weight for 1 Oxygen (15.999) = 18.015 72.74g/18.015 = 4.04 mol Mole of CaCl2 = 4.04 mol / 2 = 2.02 moles Ca(OH)2

OpenStudy (sleepyjess):

The question marks are an arrow

OpenStudy (abhisar):

Give me a sec..

OpenStudy (sleepyjess):

Ok

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (sleepyjess):

Could you check one other?

OpenStudy (abhisar):

yes...go ahead

OpenStudy (sleepyjess):

How many milliliters of magnesium metal, with a density of 1.74 g/mL, are needed to produce 102.3 grams of magnesium oxide in the single-replacement reaction below? Mg + H2O ----> MgO + H2 Atomic Mass of Mg = 24.3050 Atomic Mass of Oxygen = 15.9994 Molar Mass of MgO = 24.3050 + 15.9994 = 40.3044 g/mol 102.3g/40.3044 g/mol = 2.54 moles Mg 2.54 moles Mg x 24.3050/mole = 61.7347 g Mg V = mass / density 61.7347 g / 1.74 g/mL = 35.4797 mL

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (sleepyjess):

Thank you!

OpenStudy (abhisar):

ur most welcome :)